OpenStudy (anonymous):

please help!!! greatly appreciated! The vertical asymptote of the function y=log(x+3)-8 is a. x=-3 b. x=3 c. x=8 or d. x=-8

OpenStudy (anonymous):

this shifts the function to the left 3 units so vertical asymptote shifts from \(x=0\) to \(x=-3\)

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

and can somebody help me find the range and domain of the function y=ln(x-1)+2

OpenStudy (accessdenied):

If you think about the logarithm and where it disappears / is no longer defined, its basically when the input is equal to 0, log 0 is not defined. So we could also set x + 3 = 0. :)

OpenStudy (anonymous):

thank you so much @AccessDenied

OpenStudy (anonymous):

so did you get -3 for the answer also then?

OpenStudy (accessdenied):

Yep. :) As for that second q, the domain is once again the points where the input is greater than 0. We could simply set 'x - 1>0' here for those values. Once we hit 0, we get another asymptote and then the graph vanishes. :) In this case, the range is usually just all real numbers. It gets very very small (large/negative) as it approaches its asymptote, and it gets larger and larger as you go bigger and bigger values.

OpenStudy (anonymous):

thank you so much

OpenStudy (accessdenied):

You're welcome! :)

OpenStudy (anonymous):

so what is the domain for that question?

OpenStudy (anonymous):

i wish i could give you best answer but i have already given it for this question so i can ask another one and give you best answer there? :)

OpenStudy (accessdenied):

y = ln(x - 1) + 2 --> ln is only defined where x - 1 > 0. We just solve this inequality for the domain.

OpenStudy (anonymous):

so the domain is 1, ∞?

OpenStudy (anonymous):

or (-1,∞)

OpenStudy (accessdenied):

x - 1 > 0 add 1 to both sides: x > 1 So (1, infinity) is correct. :)

OpenStudy (anonymous):

thank you so much again!!!! xx